I need help with a data mining problem

Consider the following transaction dataset. T1: a
Answer #1

a)Support is 'the number of times {e} is in all the transactions / total number of transaction' multiplied by 100.

Support of {e} = 6/8 =3/4 = 0.75 = 75%

Support of {b,d} = 3/8 = 0.375 = 37%

Support of {b,d,e} = 2/8 = 0.25 = 25%

b)for example X->Y is an association rule the confidence is computed is as follows

Confidence is 'number of times X and Y is repeated in all the transactions / number of times X is repeated'.

therefore {b.d} -> {e} = 2/6 = 1/3 = 0.33 = 33%

{e} -> {b.d} = 2/3 = 0.66 = 66%

Confidence is not a symmetric measure

d) these are the three item set list which supports minimum support

{a,b,d} = 0.25

{b,c,e} = 0.25